# Neat: Nonlinear Parameter-efficient Adaptation of Pre-trained Models

### v2

@@ 4 Experiment
In the experiments, we evaluate the proposed Neat and answer the following questions: RQ1) How does Neat compare to state-of-the-art PEFT methods on NLP tasks? RQ2) How does Neat compare to state-of-the-art PEFT methods on vision tasks? RQ3) How does the performance of Neat vary with different fine-tuned modules, depths of the lightweight neural network, or non-linear activation functions?
Table 1: Common Reasoning performance of Neat and PEFT baselines on LLaMA 2-7B and LLaMA 3-8B. Results marked with “+” are taken from Liu et al. (2024), and those marked with $\ast$” are taken from Wang et al. (2024c). Best results are in bold. “AVG” means the average accuracy of all datasets.
Model PEFT Accuracy ($\uparrow$) BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA AVG LLaMA2-7B LoRA+ 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6 PiSSA* 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8 MiLoRA* 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2 Neat 71.9 84.0 80.4 88.9 84.6 86.5 71.6 83.0 81.4 LLaMA3-8B LoRA+ 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8 PiSSA* 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4 MiLoRA* 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9 Neat 72.1 87.0 80.9 94.3 86.7 91.4 78.9 84.8 84.5
@@ 4.1 Benchmarks and Experiment Setups
We experiment Neat on datasets from four representative benchmarks:
Commonsense Reasoning
covers diverse multi-choice problems from BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC-e and ARC-c (Clark et al., 2018), and OpenBookQA (Mihaylov et al., 2018) datasets. Following Wang et al. (2024c), we finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-8B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023) benchmark which combines all previous training sets, and evaluate the accuracy on their testing sets separately.
Arithmetic Understanding
consists of two math reasoning datasets: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-8B (AI@Meta, 2024) on MetaMath (Yu et al., 2023) dataset following Wang et al. (2024c). Models need to generate correct answers, and accuracy is used as the evaluation metric.
Natural Language Understanding
consists of eight datasets from the GLUE benchmark (Wang et al., 2018). We follow the evaluation metrics and setups from Gao et al. (2024); Wu et al. (2024a).
Image Classification
consists of Oxford-Pets (Parkhi et al., 2012), CIFAR10 (Krizhevsky, 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), RESISC45 (Cheng et al., 2017), StanfordCars (Krause et al., 2013), FGVC (Maji et al., 2013), and CIFAR100 (Krizhevsky, 2009) following Gao et al. (2024). The first five datasets have small label spaces, while the last three have large label spaces.
Baselines methods are chosen on a task basis. On each task, Neat is compared with representative baselines from the corresponding domain.
See our appendix for details about the datasets (App E), baseline methods (App C), and hyper-parameters (App D).
@@ 4.2 Performance Comparison
We showcase Neat performance on different tasks.
4.2.1 Commonsense Reasoning
@@
We experiment Neat with eight commonsense reasoning datasets to address RQ1, results are shown in Tab LABEL:tab:commonsense. We compare the performance of three state-of-the-art baselines with the proposed Neat, and Neat consistently outperforms all of them, achieving the highest accuracy on all tasks. Specifically, Neat surpasses LoRA, PiSSA, and MiLoRA in terms of average accuracy by 4.6%, 10%, and 2.5%, respectively, when using LLaMA2-7B as the backbone. On LLaMA3-8B as the backbone, Neat demonstrates average improvements of 4.9%, 11.8%, and 2.9% over LoRA, PiSSA, and MiLoRA, respectively. These results highlight the effectiveness and superiority of Neat as a PEFT method.
4.2.2 Arithmetic Reasoning
@@
In this section, we present results on two arithmetic reasoning tasks in Tab 2 to help address RQ1. From the table, while full fine-tuning (FFT) achieves highest accuracy across the two datasets, the performance gap between the proposed Neat and FFT is very small, despite that Neat relies on significantly fewer trainable parameters. Moreover, compared to state-of-the-art PEFT baselines, Neat achieves remarkable performance improvements. In terms of average accuracy, Neat demonstrates improvements of 7.5%, 12.4%, and 2.4% over LoRA, PiSSA, and MiLoRA, respectively. These results on clearly confirm that Neat is highly effective and efficient for complex reasoning tasks.
Table 2: Arithmetic Reasoning performance on LLaMA 2-7B. Results marked with “+” are taken from Yu et al. (2023), and those marked with $\ast$” are taken from Wang et al. (2024c). Best results are in bold. “AVG” means the average accuracy of all datasets.
Method GSM8K MATH AVG
FFT + 66.50 19.80 43.20
LoRA* 60.58 16.88 38.73
PiSSA* 58.23 15.84 37.04
MiLoRA* 63.53 17.76 40.65
Neat 65.05 18.30 41.68
4.2.3 Natural Language Understanding
@@
We further conduct experiments on the GLUE to answer RQ1, results are shown in Tab 3. From the table, Neat significantly outperforms state-of-the-art PEFT methods. Specifically, Neat-S, which uses a similar number of trainable parameters as FourierFT (Gao et al., 2024), DiReFT (Wu et al., 2024a), and LoReFT (Wu et al., 2024a), surpasses all PEFT baselines and experiences only a small performance drop (0.2%) compared to FFT. Additionally, Neat-L exceeds the performance of all baselines, including FFT, with roughly the same number of trainable parameters as in LoRA. These results demonstrate that Neat exhibits excellent generalization ability while maintaining great parameter efficiency.
Table 3: GLUE benchmark performance on RoBERTa-base. Results marked with $\ast$” are taken from Wu et al. (2024b). Best results are in bold. “AVG” means the average accuracy of all datasets. Neat-S applies trainable modules to layers starting from the 4th layer, with hidden dimensions set to 1. This matches the parameter numbers of FourierFT. Neat-L applies Neat to all layers with hidden dimension 8, aligning the parameter budget of LoRA.
PEFT Params (%) Accuracy ($\uparrow$) MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B AVG FFT 100% 87.3 94.4 87.9 62.4 92.5 91.7 78.3 90.6 85.6 Adapter 0.318% 87.0 93.3 88.4 60.9 92.5 90.5 76.5 90.5 85.0 LoRA 0.239% 86.6 93.9 88.7 59.7 92.6 90.4 75.3 90.3 84.7 AdapterFNN 0.239% 87.1 93.0 88.8 58.5 92.0 90.2 77.7 90.4 84.7 BitFit 0.080% 84.7 94.0 88.0 54.0 91.0 87.3 69.8 89.5 82.3 RED 0.016% 83.9 93.9 89.2 61.0 90.7 87.2 78.0 90.4 84.3 FourierFT 0.019% 84.7 94.2 90.0 63.8 92.2 88.0 79.1 90.8 85.3 DiReFT 0.015% 82.5 92.6 88.3 58.6 91.3 86.4 76.4 89.3 83.2 LoReFT 0.015% 83.1 93.4 89.2 60.4 91.2 87.4 79.0 90.0 84.2 Neat-S 0.019% 84.9 94.3 90.2 64.6 92.0 88.3 78.3 90.5 85.4 Neat-L 0.241% 86.9 95.2 90.0 64.8 92.3 90.3 82.7 90.7 86.6
@@
Table 4: Image Classification performance on ViT-base. Best results are in bold. “AVG” means the average accuracy of all datasets. Results marked with $\ast$” are taken from Gao et al. (2024).
@@
Neat 263K 93.62 80.21 98.78 79.61 98.85 52.93 94.71 92.02 86.34
4.2.4 Image Classification
In this section, we conduct experiments on image classification tasks to address RQ2, Neat uses depth of 6, and results are shown in Tab 4. From the table, Neat significantly outperforms LoRA and FourierFT using the same number of trainable parameters. Specifically, Neat achieves performance improvements of 11.05%, 7.30%, and 26.02% compared to LoRA, FourierFT, and LP, respectively. Furthermore, compared to FFT, the proposed Neat shows negligible performance drop (86.49% v.s. 86.34%), while using only 0.3% of the trainable parameters required by FFT. This demonstrates that Neat exhibits exceptional adaptation capability not only on NLP tasks, but also on vision tasks as well. Additionally, it verifies the effectiveness of the nonlinear adaptation used in Neat.
Table 5: Ablation Study on image classification task. The parameters count is the same and “AVG” means the average accuracy of all datasets. For simple and fair comparison, Neat uses depth of 2.
@@ 4.2.4 Image Classification
Neat 93.77 80.03 98.70 77.57 98.79 53.60 94.27 92.47 86.15
4.3 Ablation Study
In this section, we present an ablation study with two variants of LoRA to validate the effectiveness of our proposed framework: 1) nonlinear LoRA $\bm{y}=(\bm{W_{0}+\sigma(\mathbf{A})\mathbf{B}})\bm{x}$, and 2) multiplicative LoRA $\bm{y}=(\mathbf{W}_{0}+\mathbf{W}_{0}\mathbf{A}\mathbf{B})\bm{x}$. Experiments are conducted on image classification benchmarks, and results are reported in Tab 5. According to the table, both nonlinear LoRA and multiplicative LoRA perform worse than Neat. This highlights the effectiveness of incorporating nonlinear approximations and explicitly using model weights as input to the nonlinear function in Neat.
@@ 4.3 Ablation Study
The expressiveness of neural networks grow with its depth (Raghu et al., 2017). This implies that deeper Neat may be able to capturing more complex patterns required by ideal weight updates. In this section, we experiments Neat with varying depth on vision, NLU, commonsense reasoning, and arithmetic reasoning benchmarks to address RQ3. Results are shown in Fig 2. See Neat layer architecture in Fig 5 and more details and results in App F To ensure fair comparison, we fix all other hyper-parameters when varying depth.
From Fig 2, increasing network depth leads to better performance. Specifically, when using depth of 6, the classification accuracy achieves 81.04%, marking a 1.7% improvement compared with using depth of 2. On RTE, using 6 layers also boosts the performance by a remarkable $4\%$ compared to using 2. In addition, deeper networks tend to bring more improvement on complex benchmarks, such as MATH and PIQA. Finally, we would like to highlight that Neat intermediate layers consist much fewer parameters ($\mathbb{R}^{r\times r}$, where $r$ is the hidden layer dimension) compared to the pre-trained model’s weight. Therefore, the additional parameter overhead of stacking more hidden layers is negligible and does not affect the parameter efficiency of Neat. These results further validate the effectiveness of introducing non-linear adaptation.
Figure 2: Accuracy on the RTE, StanfordCars, PIQA and MATH dataset with varying depths of the neural network used in Neat. The depth here represents the total number of layers in the neural network. We choose depth equals to 2, 4 and 6 layers in the figure.
Figure 3: Influence of different nonlinear activations choices for NEAT. Experiments are conducted on StanfordCars, Neat depth is fixed to 2. Different activations share a similar pattern of dependency on learning rate.
@@
One key innovation of Neat compared to LoRA and other PEFT methods, which rely solely on linear transformations for modeling weight updates, is the introduction of non-linear activations within the adaptation neural network. Since the choice of non-linear activations directly affects the learning process and the dynamics of weight updates, we investigate how different non-linear activations affects the adaptation performance to address RQ3. To this end, we perform experiments on the StanfordCars benchmark using various non-linear activations, including ReLU, Leaky ReLU, GELU, Tanh, and sinusoidal activation ($\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$). Corresponding results are presented in Fig 3. To ensure a fair comparison, the number of trainable parameters is fixed. We optimize other hyperparameters such as learning rate for better performance.
From the figure, the best performance achieved by different activation functions is similar, indicating that the adaptation potential of various activations is comparable. This implies that Neat can benefit from various type of nonlinearity induced by different activations. However, it is also worth noting that sinusoidal activations encounters a performance drop at large learning rates. Consequently, tuning basic hyperparameters such as learning rate can still be beneficial. In conclusion, we suggest ReLU as a default choice in execution, given its practical simplicity (Teney et al., 2024).
@@
Figure 4: Accuracy of Neat with different targeted fine-tuning modules, including just QV layers and a combination of QV and MLP layers, on image classification datasets.
We end up this section with a study on applying Neat to different modules in a ViT, to help better understand RQ3. Specifically, given the importance of MLP in Transformer architecture, we compare two settings: 1) Following Hu et al. (2021a), we apply Neat to the query and value layers (QV layers) in the multi-head self-attention module (MHSA) in ViT. 2) Besides QV layers, we also apply Neat to MLP layers. We tune the hidden dimension $r$ to ensure the same parameter scale for fair comparison, and tune the hyperparameters to maximize performance. Corresponding results are shown in Fig. 4.
From the figure, applying Neat to the QV layers yields results comparable to applying Neat to both the QV and MLP layers. This indicates that Neat is robust to the selections of fine-tuning different modules. This finding confirms another key advantage of Neat: it does not require extensive manual tuning on which parts (modules, layers) of the foundation model Neat should be applied. Consequently, Neat can be easily incorporated to a wide range of scenarios.

### v3

@@ 5 Experiment
In the experiments, we evaluate the proposed PEANuT and answer the following questions:
  1. RQ1
    How does PEANuT compare to state-of-the-art PEFT methods on NLP and vision tasks?
  2. RQ2
    What is the role of nonlinear approximation in the proposed PEANuT?
  3. RQ3
    What is the real runtime and memory consumption of proposed PEANuT?
  4. RQ4
    How does the performance of PEANuT vary with different fine-tuned modules, depths of the lightweight neural network, or non-linear activation functions?
@@ 5.1 Benchmarks and Experiment Setups
We experiment PEANuT on datasets from four representative benchmarks: 1) Commonsense Reasoning covers diverse multi-choice problems from BoolQ (boolq), PIQA (piqa), SIQA (siqa), HellaSwag (hellaswag), WinoGrande (winogrande), ARC-e and ARC-c (ARC), and OpenBookQA (openbookqa) datasets. Following milora, we finetune LLaMA2-7B (llama2-7b), LLaMA3-8B (llama3) and Qwen3-8B (qwen3) on Commonsense170K (llmadapter) benchmark which combines all previous training sets, and evaluate the accuracy on their testing sets separately. 2) Arithmetic Understanding consists of two math reasoning datasets: GSM8K (gsm8k) and MATH (MATH). We finetune LLaMA2-7B (llama2-7b) and Qwen3-8B (qwen3) on MetaMath (metamath) dataset following milora. Models need to generate correct answers, and accuracy is used as the evaluation metric. 3) Natural Language Understanding consists of eight datasets from the GLUE benchmark (glue). We follow the evaluation metrics and setups from fourierft, wuandarora2024reft. 4) Image Classification consists of Oxford-Pets (pets), CIFAR10 (cifar), DTD (dtd), EuroSAT (euro), RESISC45 (resisc), StanfordCars (cars), FGVC (fgvc), and CIFAR100 (cifar) following fourierft. The first five datasets have small label spaces, while the last three have large label spaces.
Baselines methods are constructed on a task basis. Specifically, for each task, the proposed PEANuT is compared with representative baselines from the corresponding domain. For both Commonsense Reasoning and Arithmetic Understanding, following milora, LoRA (lora), PiSSA (pissa) and MiLoRA (milora) are employed as baselines. PEANuT is applied to query, key, value, MLP up and MLP down layers. For Natural Language Understanding, we follow the setup from prior works (fourierft, wuandarora2024reft) that evaluate various representative PEFT methods, including LoRA (lora), Adapter houlsby2019parameter, BitFit (zaken2021bitfit), RED (wu2024advancing), DoRA (liu2024dora), ReFT wuandarora2024reft, and FourierFT (fourierft). For Image Classification, we follow the setting of fourierft and take linear probing (LP), LoRA (lora) and FourierFT (fourierft) as baselines. PEANuT is applied to the query and value layers. See our appendix for details about the datasets (App D) and hyper-parameters (App C).
@@ 5.2 Performance Comparison
We showcase PEANuT performance on different tasks.
@@
Commonsense Reasoning. We experiment PEANuT with eight commonsense reasoning datasets to address RQ1, results are shown in Tab LABEL:tab:commonsense. We compare the performance of three state-of-the-art baselines with the proposed PEANuT, and PEANuT consistently outperforms all of them, achieving the highest accuracy on all tasks. Specifically, PEANuT surpasses LoRA, PiSSA, and MiLoRA in terms of average accuracy by 4.6%, 10%, and 2.5%, respectively, when using LLaMA2-7B as the backbone. On LLaMA3-8B as the backbone, PEANuT demonstrates average improvements of 4.9%, 11.8%, and 2.9% over LoRA, PiSSA, and MiLoRA, respectively. With Qwen3-8B as the backbone, PEANuT improves average accuracy over LoRA and MiLoRA by 3.9% and 2.4%, respectively. These results highlight the effectiveness and superiority of PEANuT as a PEFT method.
@@
Arithmetic Reasoning. In this section, we present results on two arithmetic reasoning tasks in Tab 4 to help address RQ1. From the table, while full fine-tuning (FFT) achieves highest accuracy across the two datasets, the performance gap between the proposed PEANuT and FFT is very small, despite that PEANuT relies on significantly fewer trainable parameters. Moreover, compared to state-of-the-art PEFT baselines, PEANuT achieves remarkable performance improvements. In terms of average accuracy, PEANuT demonstrates improvements of 7.5%, 12.4%, and 2.4% over LoRA, PiSSA, and MiLoRA, respectively, when using LLaMA2-7B as the backbone. With Qwen3-8B as the backbone, PEANuT improves average accuracy over LoRA and MiLoRA by 7.1% and 3.7%. These results on clearly confirm that PEANuT is highly effective and efficient for complex reasoning tasks.
@@
Natural Language Understanding. We further conduct experiments on the GLUE to answer RQ1, results are shown in Tab 3. From the table, PEANuT significantly outperforms state-of-the-art PEFT methods. Specifically, PEANuT-S, which uses a similar number of trainable parameters as FourierFT (fourierft), DiReFT (wuandarora2024reft), and LoReFT (wuandarora2024reft), surpasses all PEFT baselines and experiences only a small performance drop (0.2%) compared to FFT. Additionally, PEANuT-L exceeds the performance of all baselines, including FFT, with roughly the same number of trainable parameters as in LoRA. These results demonstrate that PEANuT exhibits excellent generalization ability while maintaining great parameter efficiency.
Image Classification. In this section, we conduct experiments on image classification tasks to address RQ2, PEANuT uses depth of 6, and results are shown in Tab 2. From the table, PEANuT significantly outperforms LoRA and FourierFT using the same number of trainable parameters. Specifically, PEANuT achieves performance improvements of 11.05%, 7.30%, and 26.02% compared to LoRA, FourierFT, and LP, respectively. Furthermore, compared to FFT, the proposed PEANuT shows negligible performance drop (86.49% v.s. 86.34%), while using only 0.3% of the trainable parameters required by FFT. This demonstrates that PEANuT exhibits exceptional adaptation capability not only on NLP tasks, but also on vision tasks as well. Additionally, it verifies the effectiveness of the nonlinear adaptation used in PEANuT.
@@
Table 1: Common Reasoning performance of PEANuT and PEFT baselines on LLaMA 2-7B, LLaMA 3-8B and Qwen 3-8B. Results marked with “+” are taken from liu2024dora, and those marked with $\ast$” are taken from milora. Best results are in bold. “AVG” means the average accuracy of all datasets.
Model PEFT Accuracy ($\uparrow$) BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA AVG LLaMA2-7B LoRA+ 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6 PiSSA* 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8 MiLoRA* 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2 PEANuT 71.9 84.0 80.4 88.9 84.6 86.5 71.6 83.0 81.4 LLaMA3-8B LoRA+ 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8 PiSSA* 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4 MiLoRA* 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9 PEANuT 72.1 87.0 80.9 94.3 86.7 91.4 78.9 84.8 84.5 Qwen3-8B LoRA 86.3 87.2 84.1 92.5 81.5 89.6 78.8 89.5 86.2 MiLoRA 85.2 89.3 84.2 94.6 82.2 92.3 82.7 89.5 87.5 PEANuT 89.4 90.2 87.4 95.7 85.5 92.7 82.6 93.6 89.6
Table 2: Image Classification performance on ViT-base. Best results are in bold. “AVG” means the average accuracy of all datasets. Results marked with $\ast$” are taken from fourierft.
@@
PEANuT 263K 93.62 80.21 98.78 79.61 98.85 52.93 94.71 92.02 86.34
Table 3: GLUE benchmark performance on RoBERTa-base. Results marked with $\ast$” are taken from wu2024advancing. Best results are in bold. “AVG” means the average accuracy of all datasets. PEANuT-S applies trainable modules to layers starting from the 4th layer, with hidden dimensions set to 1. This matches the parameter numbers of FourierFT. PEANuT-L applies PEANuT to all layers with hidden dimension 8, aligning the parameter budget of LoRA.
PEFT Params (%) Accuracy ($\uparrow$) MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B AVG FFT 100% 87.3 94.4 87.9 62.4 92.5 91.7 78.3 90.6 85.6 Adapter 0.318% 87.0 93.3 88.4 60.9 92.5 90.5 76.5 90.5 85.0 LoRA 0.239% 86.6 93.9 88.7 59.7 92.6 90.4 75.3 90.3 84.7 AdapterFNN 0.239% 87.1 93.0 88.8 58.5 92.0 90.2 77.7 90.4 84.7 BitFit 0.080% 84.7 94.0 88.0 54.0 91.0 87.3 69.8 89.5 82.3 RED 0.016% 83.9 93.9 89.2 61.0 90.7 87.2 78.0 90.4 84.3 FourierFT 0.019% 84.7 94.2 90.0 63.8 92.2 88.0 79.1 90.8 85.3 DiReFT 0.015% 82.5 92.6 88.3 58.6 91.3 86.4 76.4 89.3 83.2 LoReFT 0.015% 83.1 93.4 89.2 60.4 91.2 87.4 79.0 90.0 84.2 PEANuT-S 0.019% 84.9 94.3 90.2 64.6 92.0 88.3 78.3 90.5 85.4 PEANuT-L 0.241% 86.9 95.2 90.0 64.8 92.3 90.3 82.7 90.7 86.6
Table 4: Arithmetic Reasoning performance on LLaMA 2-7B and Qwen 3-8B. Results marked with “+” are taken from metamath, and those marked with $\ast$” are taken from milora. Best results are in bold. “AVG” means the average accuracy of all datasets.
Model Method GSM8K MATH AVG
LLaMA2-7B FFT + 66.50 19.80 43.20
LoRA* 60.58 16.88 38.73
PiSSA* 58.23 15.84 37.04
MiLoRA* 63.53 17.76 40.65
PEANuT 65.05 18.30 41.68
Qwen3-8B LoRA 85.22 67.26 76.24
MiLoRA 89.01 68.58 78.80
PEANuT 92.87 70.50 81.69
5.3 Ablation Study
In this section, in order to answer RQ2, we present an ablation study with two variants of LoRA to validate the effectiveness of our proposed framework: 1) nonlinear LoRA $\bm{y}=(\bm{W_{0}+\sigma(\mathbf{A})\mathbf{B}})\bm{x}$, and 2) multiplicative LoRA $\bm{y}=(\mathbf{W}_{0}+\mathbf{W}_{0}\mathbf{A}\mathbf{B})\bm{x}$. Experiments are conducted on image classification benchmarks, and results are reported in Tab 5. According to the table, both nonlinear LoRA and multiplicative LoRA perform worse than PEANuT. This highlights the effectiveness of incorporating nonlinear approximations and explicitly using model weights as input to the nonlinear function in PEANuT.
Table 5: Ablation Study on image classification task. The parameters count is the same and “AVG” means the average accuracy of all datasets. For simple and fair comparison, PEANuT uses depth of 2.
@@ 5.3 Ablation Study
PEANuT 93.77 80.03 98.70 77.57 98.79 53.60 94.27 92.47 86.15
Table 6: Accuracy comparison of PEANuT using RoBERTa-base with different depth configurations on the GLUE benchmark. The highest accuracy of methods per category are in bold. “AVG” means the average accuracy of all datasets.
Figure 2: Implementation of introducing more depths to PEANuTt. We insert multiple intermediate layers into the layers from vanilla PEANuT, with non-linear activation in between. The depth is described as the number of layers in PEANuT, with vanilla PEANuT having a depth of 2 (i.e. the input and output layers).
Table 7: Runtime and memory consumption of proposed PEANuT.
Dataset Method Time Memory
MRPC LoRA 77.7s 6916MB
PEANuT 78.4s 6916MB
SST-2 LoRA 870.7s 2410MB
PEANuT 911.4s 2410MB
Commonsense
Reasoning
LoRA 5.6h 22.7GB
PEANuT 5.7h 23.8GB
5.4 Runtime and Memory Cost
To answer RQ3, we evaluate the computational efficiency of our proposed method, PEANuT, by measuring its runtime and memory consumption across three representative datasets: MRPC, SST-2, and Commonsense Reasoning. Table 7 summarizes the results, comparing PEANuT against the LoRA approach under identical settings. For a fair comparison, we ensure that the number of trainable parameters is matched between PEANuT and LoRA. All experiments are conducted on the same hardware setup using a single NVIDIA A100 GPU. As shown in Table 7, PEANuT exhibits comparable runtime and memory usage to LoRA across all tasks. On the MRPC and SST-2 datasets, PEANuT incurs only marginal overhead in training time, with identical memory consumption. For the larger Commonsense Reasoning dataset, PEANuT takes 5.7 hours and 23.8GB of memory, compared to LoRA’s 5.6 hours and 22.7GB. The slightly higher memory usage is attributed to the additional nonlinear transformation module in PEANuT, but the increase remains slight. Overall, the results demonstrate that PEANuT achieves improved performance (as discussed in earlier sections) with negligible additional cost in training runtime and memory, highlighting its practicality and scalability for real-world deployment.
Figure 3: Accuracy on the RTE, StanfordCars, PIQA and MATH dataset with varying depths of the neural network used in PEANuT. The depth here represents the total number of layers in the neural network. We choose depth equals to 2, 4 and 6 layers in the figure.
Figure 4: Influence of different nonlinear activations choices for PEANuT. Experiments are conducted on StanfordCars, PEANuT depth is fixed to 2. Different activations share a similar pattern of dependency on learning rate.
Figure 5: Accuracy of PEANuT with different targeted fine-tuning modules, including just QV layers and a combination of QV and MLP layers, on image classification datasets.
@@ 5.4 Runtime and Memory Cost
To answer RQ4, we analyze the impact of depth on the performance of PEANuT. Deeper architectures are generally more expressive and can better model the complex, nonlinear relationships involved in ideal weight updates (raghu2017expressive). We evaluate PEANuT with varying depth across NLU, vision, commonsense reasoning, and arithmetic reasoning tasks.
We increase the number of intermediate layers inserted between PEANuT’s input and output projections. Each intermediate layer is a small feedforward block of shape $\mathbb{R}^{r\times r}$ with non-linear activations. These layers are lightweight compared to the input/output projections ($\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times d_{2}}$), and add minimal overhead since $r\ll d_{2}$. The adaptation starts from the frozen base weight $\mathbf{W}^{0}$, which is transformed through multiple layers to predict $\Delta\mathbf{W}$. We adopt residual connections for stable optimization and improved convergence. All other hyperparameters are kept fixed during this analysis. The layer structure is illustrated in Fig. 2.
The results in Table 6 (GLUE) and Fig. 3 (RTE, Cars, PIQA, MATH) indicate that increasing depth consistently improves accuracy. For instance, average GLUE accuracy increases from 86.0 to 86.6 when moving from 2 to 6 layers, with no significant change in parameter count. On other benchmarks, deeper configurations yield steady gains up to 6 layers. Beyond this, performance may slightly drop (e.g., at depth 10), likely due to optimization difficulties without fine-grained hyperparameter tuning.
In summary, depth enhances PEANuT’s effectiveness across tasks, offering better adaptation capability with negligible cost in memory or parameters. However, very deep settings may require further tuning to maintain stability.
@@
One key innovation of PEANuT compared to LoRA and other PEFT methods, which rely solely on linear transformations for modeling weight updates, is the introduction of non-linear activations within the adaptation neural network. Since the choice of non-linear activations directly affects the learning process and the dynamics of weight updates, we investigate how different non-linear activations affects the adaptation performance to address RQ4. To this end, we perform experiments on the StanfordCars benchmark using various non-linear activations, including ReLU, Leaky ReLU, GELU, Tanh, and sinusoidal activation ($\sigma_{\textnormal{p}}(x)=\sin(2\pi x)$). Corresponding results are presented in Fig 4. To ensure a fair comparison, the number of trainable parameters is fixed. We optimize other hyperparameters such as learning rate for better performance.
From the figure, the best performance achieved by different activation functions is similar, indicating that the adaptation potential of various activations is comparable. This implies that PEANuT can benefit from various type of nonlinearity induced by different activations. However, it is also worth noting that sinusoidal activations encounters a performance drop at large learning rates. Consequently, tuning basic hyperparameters such as learning rate can still be beneficial. In conclusion, we suggest ReLU as a default choice in execution, given its practical simplicity (teney2024neuralredshift).
@@
We end up this section with a study on applying PEANuT to different modules in a ViT, to help better understand RQ4.
Specifically, given the importance of MLP in Transformer architecture, we compare two settings: 1) Following lora, we apply PEANuT to the query and value layers (QV layers) in the multi-head self-attention module (MHSA) in ViT. 2) Besides QV layers, we also apply PEANuT to MLP layers. We tune the hidden dimension $r$ to ensure the same parameter scale for fair comparison, and tune the hyperparameters to maximize performance. Corresponding results are shown in Fig. 5.
From the figure, applying PEANuT to the QV layers yields results comparable to applying PEANuT to both the QV and MLP layers. This indicates that PEANuT is robust to the selections of fine-tuning different modules. This finding confirms another key advantage of PEANuT: it does not require extensive manual tuning on which parts (modules, layers) of the foundation model PEANuT should be applied. Consequently, PEANuT can be easily incorporated to a wide range of scenarios.